OpenStudy (anonymous):

what are the common denominators between 25 and 20

OpenStudy (solomonzelman):

\[hint:~~~~~~~~20= 4 \times 5~~~~~~~~~~~~~~25=5 \times 5 \] example! common denominator btw 6 and 8 \[6 = 2 \times 3\]\[8=2 \times 4\]\[common~~denominator=~~2 \times 3 \times 4=24\]

OpenStudy (solomonzelman):

(b/c they in my ex. have 2 in common)

OpenStudy (solomonzelman):

tel me what you get, to verify that you are correct ( please ) .

OpenStudy (anonymous):

i still don't understand if the problem is 11/25 and 9/20 i have to find the common denominator

OpenStudy (solomonzelman):

yes, if you want to add or subtract them.

OpenStudy (solomonzelman):

what exactly do you need to do ?